Wednesday, August 9, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. Latif SLides
  2. Commissioning plan updates from detector systems
  3. Status update from detector owners, calibrations, efficiencies, etc.
  4. Review systems owners and update information
  5. Address system specific points of interest for the Fall 2017 run

Minutes

General

  • Brad would like to add a few items to the wish list
    • We need spare VXS crate
    • Intel single board computers
  • Simona has coincidence modules in the hall that she required for her cosmic test stand setup
    • These were taken from the Hall-C counting house

Trigger/DAQ

  • GEM crate integration is now an issue see logbook entry:
    • Most likely a TIpcie issue which has been observed in the past
  • New fiber optic pull for the HMS has been done
    • FE group is working on completing this
    • Latencies will change since there is one less path for the SHMS
    • Once everything is integrated new latencies in the CRLs will need to be set
  • Eric and Carlos will begin conducting work on the production trigger
    • If folks require the DAQ please contact Eric so that he can ensure the DAQ is in a stable state
  • Steve reported that he is working on the TI Scalers & proper gating of regular scalers

    • He will take the HMS DAQ for a while
    • Log book entry will follow to alert folks

HMS

12 GeV Drift Chambers

  • A voltage scan of one chambers has been conducted
    • A few issues have been identified
  • Some of the efficiencies of the planes appear to be lower than expected
    • This is a work in progress
    • A hot wire has been located
  • DC1 is nearly complete
  • DC2 will be tested immediately thereafter
  • The cabeling for all channels is ready
  • The aim is too have DC2 wrapped up by the end of the month
  • They have to mounted in ESB and fiducialized and then moved to hall

SHMS

GEM

  • Latif showed some slides
  • There exists a git repo for development
  • Readme contains instructions of how to incorporate the GEM analyzer into the hall analyzer
  • The class is THcGEM which is traded in a similar fashion to the other detector classes
  • The GEM data has been shown to exist in the Tree along with other detector associated variables
  • The mapping still needs to be corrected
  • The data that was shown is the data that was taken during latency scans a few weeks ago
  • There is a clear plan moving forward

Drift Chambers

  • The chamber seems to be gas tight
  • Brad is working with Chuck L. to verify that the chamber can hold HV while filled with nitrogen
  • If functional it will be moved to the ESB for further testing

Hodoscopes

  • The PMTs have been removed from 17 bars
  • 13 bars have been cleaned and are ready to be coupled to PMT’s
  • The cookies were in fact thicker than they should have been
  • Some of the PMTs are being delivered to Carl Z. For QE testing
  • Some cookies were 500 um while some were < 200 um
  • There exists a correlation between thick cookies and the XP PMT’s as opposed to the ET’s
  • The new cookie fabrication is proceeding nicely
    • Sometimes they sheer during fabrication although
    • A solution is being worked out
  • The new tubes should be shipped from the UK on Aug. 17
    • This is okay due to the accelerator delay
  • Mid September is the nominal goal to have the plane assembled and being tested in the cosmic test stand